OpenStudy (anonymous):

Instructions:Type the correct answer in each box. Use numerals instead of words. If necessary, use / for the fraction bar(s). Find the missing term and the missing coefficient. (6a −______)5a =______a^2 − 35a

OpenStudy (anonymous):

Some help would be appreciated greatly...

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

5 * ? = -35

OpenStudy (igreen):

That's the same as: 5x = -35 Divide 5 to both sides, what do you get?

OpenStudy (anonymous):

-7 but how did you get to that point?

OpenStudy (igreen):

See the -35a? We want to multiply 5a to something to get that, and that number is -7. -7 * 5a = -35a

OpenStudy (igreen):

So 7 goes in the first blank.

OpenStudy (igreen):

Now, you see 6a? We multiply 5a * 6a, what do you get?

OpenStudy (anonymous):

30a?

OpenStudy (igreen):

Close, but remember we also multiply the two variables together. \(a \times a = a^2\) So it's actually 30a^2. So 30 goes in the 2nd blank.

OpenStudy (anonymous):

So how did you solve the problem in the beginning. It's just been quite a while for me. I'm sorry.

OpenStudy (igreen):

\((6a -~?)5a =~?a^2 - 35a\) Okay, we're supposed to distribute 5a into the parenthesis, to do that we multiply it to every term in the parenthesis. So we multiply 5a * 6a, which gives us 30a^2. And it also tells us that the result of distributing gives us -35a So we find out 5a * what = -35a, which is -7.

OpenStudy (igreen):

Understand?

OpenStudy (anonymous):

Thank you, yes.